In a population of wildflowers, flower color is controlled by a single locus. Homozygous dominant individuals are red, homozygous recessive individuals are white, and heterozygous individuals are pink. If 16 of 100 flowers are white and 32 of 100 are pink, can you calculate the expected value for FST?
Answer #1

RR (red) = 52

Rr (pink) = 32

rr (white) = 16

Calculate allele frequency

Each homozygote will have two alleles, each heterozygote will have one allele. Note that the denominator will be twice Ni (twice as many alleles as individuals)

P = 2*52 + 32/200 = 0.68

q = 1- p = 0.32

Calculate the local observed heterozygosity

Hobs = 32/100 = 0.32

H exp = 1 - (p2 + q2) = 1 - (0.682 + 0.322 ) = 1 - 0.5648 = 0.4352

Fst = Hexp - Hobs / Hexp = 0.4352 - 0.32 / 0.4352

The expected values for FST = 0.264